Let K be a field and (c, d) a pair of relatively prime positive integers. We explain the computation of the Grothendieck group of algebras of the type K[z, x_{1}, x_{2}, ...x_{n}], with relation z^{c} = f(x_{1}, x_{2}...x_{n}) where f is isobaric of weight d.